In a pile of coins, there are 15 more quarters than loonies($1) the total amount of quarters and loonies is $21.25 how many quarters are there?
Answer by fractalier(6550)

Call the number of quarters and loonies, x and y.
Then we know x = y + 15, or x - y = 15.
The value equation is
.25x+(1)y = 21.25
We can add the first equation from the second and get
.25x + y = 21.25
x - y = 15
and
1.25x=36.25
and
x= 29 quarters